A student wants to measure the density of the unknown material using the displacement method.

He put 8.4 g of unknown material into the water initially at 45.4 ml, and the water level rose to 48.5 mL. What should be the correct density (g/cm3)of the unknown material from the following? 


1
Expert's answer
2021-05-26T07:12:42-0400

Weight of the unknown = 8.4 gm

Volume of water initially "(V_1)=45.5 ml"

Final volume of water "(V_2)=48.5 ml"

So the volume of the unknown material "=V_2-V_1=48.5-45.5=3ml"


Now density:


"density=\dfrac{mass }{volume}=\dfrac{8.4}{3}=2.8 g/ml" OR "2.8g\/cm^3"